an architect designs a rectangular flower bed, such as the width is exactly two thirds the length, if the total of the fence to be used to enclose the garden is 300 feet find the diminsions
Answer by jorel1380(3719)

2(l)+2(2/3l)=300
2l+4l/3=300
6l+4l=900
10l=900
l=90
w=2/3l=60
The flower bed needed is 90' by 60'..